OpenStudy (anonymous):

how would i simplify 4sqrt125

OpenStudy (mertsj):

\[4\sqrt{125}=4\sqrt{25(5)}\]

OpenStudy (mertsj):

Does that help?

OpenStudy (anonymous):

not really..

OpenStudy (mertsj):

What is the square root of 25?

OpenStudy (mertsj):

\[4\sqrt{25(5)}=4\sqrt{25}\sqrt{5}=4(5)\sqrt{5}=20\sqrt{5}\]